Circular No. 8228

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003ir, 2003it, 2003iu, 2003iv, AND 2003iw T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report on spectra (range 370–750 nm) of several recently discovered supernovae, with all the spectrograms obtained by M. Calkins at the Mt. Hopkins 1.5-m telescope (+ FAST). The following type-Ia supernovae are listed with the time the spectrogram was obtained, the supernova spectral-feature age (given in days prior to maximum light), and their expansion velocities for Si II (rest 635.5 nm), with the adopted NASA/IPAC Extragalactic Database (NED) recession velocity for the host galaxies given parenthetically: SN 2003it (cf. IAUC 8225), Oct. 19.18 UT, 4 ± 2, 11900 km/s (7531 km/s); SN 2003iu (cf. IAUC 8226), Oct. 20.38, 0 ± 2, 10400 (12320); SN 2003iv (cf. IAUC 8226), Oct. 19.38, 2 ± 2, 11100 (10285). A spectrogram of SN 2003ir (cf. IAUC 8221), obtained on Oct. 20.50, shows it to be a type-II supernova; adopting the NED recession velocity of 7657 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 5700 km/s. A spectrogram of SN 2003iw (cf. IAUC 8226), obtained on Oct. 20.14, shows a blue, almost featureless continuum; there is a weak feature that may be an incipient P-Cyg line of Hα, implying that this is an early type-II supernova.

SUPERNOVA 2003id IN NGC 895 M. Hamuy and M. Roth, Carnegie Observatories, report that a spec-

trogram (range 380–930 nm) of SN 2003id (cf. IAUC 8201, 8203), obtained on Oct. 16.27 UT with the Dupont 2.5-m telescope (+ WFCCD) at Las Campanas, reveals unusual features. While the spectrum is characterized by strong absorptions between 466 and 536 nm (presumably due to Fe II blends) that are seen in other type-Ic events, a strong unidentified absorption centered at 562 nm distinguishes it from other type-Ic or type-Ib supernovae. This feature is blended with another strong absorption centered at 579 nm. The spectrum shows a strong Ca II triplet line exhibiting a broad P-Cyg profile; adopting the NED host-galaxy recession velocity of 2288 km/s, the minimum of the Ca II triplet at 826 nm implies an expansion velocity of −13000 km/s, which is very high for a supernova that is at least one month old. The optical lightcurves are also unusual: the V -band flux decreased by ∼ 0.25 mag between Sept. 19 and 25; a plateau phase (V = 17.25) ensued through Oct. 3; and the flux has been dropping afterwards, reaching V = 18.6 on Oct. 16. Further observations at optical and other wavelengths would be useful.

2003 October 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green